### v2

@@ 1 Introduction
A natural question that arises in the study of identifying codes (like for the usual dominating sets) is the one of extremal values for the identification number: how large can it be, with respect to some relevant graph parameters? When only the order $n$ of the graph is considered, it is known that the identification number of an identifiable graph with at least one edge lies between $\log_{2}(n+1)$ [33] and $n-1$ [23]; both values are tight and the extremal examples have been characterized in [36] and [15], respectively.
@@
With respect to the identification number, it was observed in [15] that when the maximum degree $\Delta$ of the graph $G$ is small enough with respect to the order $n$ of the graph, the ($n-1$)-upper bound can be significantly improved (for connected graphs) to $n-\frac{n}{\Theta(\Delta^{5})}$. The latter was thereafter subsequently reduced to $n-\frac{n}{\Theta(\Delta^{3})}$ in [18]. This raises the question of what is the largest possible identification number of a connected identifiable graph of order $n$ and maximum degree $\Delta$. Towards this problem, the following conjecture was posed; the goal of this paper is to study this conjecture.
@@
Even paths 2 $1$ [5] (Theorem 5)
@@ 2.1 Definitions and notations
For any vertex $v$ of $G$, the symbol $\deg_{G}(v)$ denoting the degree of the vertex $v$ in $G$ is the total number of neighbors of $v$ in $G$. A leaf of a graph $G$ is a vertex of degree $1$ in $G$. In the literature, a leaf is also known as a pendant vertex. The (only) neighbor of a leaf $v$ in a graph $G$ is called the support vertex of $v$ in $G$. The number of leaves and support vertices in $G$ are denoted by $\ell(G)$ and $s(G)$, respectively. Naturally, any vertex of a graph $G$ that is not a leaf of $G$ is usually referred to as a non-leaf vertex of $G$. The length (or the number of edges) of a longest induced path in a graph $G$ is called the diameter of $G$ and is denoted by the symbol ${\rm diam}(G)$.
On many occasions throughout this article, we shall have the need to look at a subgraph of a graph $G$ formed by deleting away some vertices or edges from $G$. To that end, given a graph $G$ and a set $S$ containing some vertices and edges of $G$, we define $G-S$ as the subgraph of $G$ obtained by deleting from $G$ all vertices (and edges incident with them) and edges of $G$ in $S$. For any positive integer $p$, let $[p]$ denote the set $\{1,2,\ldots,p\}$ and for any two distinct integers $p$ and $q$ with $p<q$, let $[p,q]$ denote the set $\{p,p+1,\ldots,q\}$.
@@ 2.2 Paths
$\gamma^{{\rm ID}}(P_{n})=\begin{cases}\frac{n}{2}+\frac{1}{2},&\text{if $n\geq 1$ is odd},\vspace{0.1cm}\\ \frac{n}{2}+1,&\text{if $n\geq 4$ is even}.\end{cases}$
@@ 3 An upper bound relating the identification number of a tree with its domination number
Let us first assume that $s$ is adjacent to $t\geq 2$ leaves $l_{1},l_{2},\dots,l_{t}$. Consider the tree $T_{s}=T-\{s,l_{1},l_{2},\dots,l_{t}\}$. Since $T_{s}$ has $n-t-1$ vertices, by induction we have $\gamma^{{\rm ID}}(T_{s})\leq n-t-1-\gamma(T_{s})$ unless $T_{s}=P_{4}$ or $T_{s}$ contains at most two vertices. However, if $T_{s}=P_{4}$, then $\gamma(T)=2=s(T)$ if vertex $s$ is adjacent to a leaf of $T_{s}$, and if $s$ is adjacent to a support vertex of $T_{s}$, then $\gamma(T)=3=s(T)$. Since $\gamma^{{\rm ID}}(T)\leq n-s(T)$ by Lemma 6, the claim holds in these cases. Furthermore, if $T_{s}$ contains one or two vertices, then again $\gamma(T)=s(T)$ and the claim follows. Therefore, we may assume by induction that there exists an identifying code $C_{s}$ in $T_{s}$ which has cardinality at most $n-t-1-\gamma(T_{s})$, with $D_{s}$ a minimum-size dominating set in $T_{s}$. Furthermore, set $C_{s}\cup\{l_{1},\dots,l_{t}\}$ is an identifying code of $T$, and set $D_{s}\cup\{s\}$ is a dominating set of $T$. Thus, $\gamma^{{\rm ID}}(T)\leq n-t-1-\gamma(T_{s})+t\leq n-\gamma(T)$, as claimed.
Let us assume next that support vertex $s$ has degree 2, and denote by $u$ the non-leaf adjacent to $s$ (possibly, $u=x$) and by $l$ the leaf adjacent to $s$. If vertex $u$ also has degree 2, then we consider the tree $T_{u}=T-\{l,s,u\}$. If $T_{u}$ contains two or fewer vertices, then $T$ is a path, a contradiction. Moreover, if $T_{u}\cong P_{4}$, then $T$ is path if $u$ is adjacent to a leaf of $T_{u}$, again a contradiction. Hence, $T$ has three support vertices and $\gamma^{{\rm ID}}(T)\leq n-3$ (consider the identifying code formed by all non-leaf vertices of $T$) while $\gamma(T)=3$. Hence, by induction, we may assume that $T_{u}$ has an optimal identifying code $C_{u}$ with cardinality $|C_{u}|=\gamma^{{\rm ID}}(T_{u})\leq n-3-\gamma(T_{u})$ and we consider a minimum-size dominating set $D_{u}$ of $T_{u}$. Notice that $D_{u}\cup\{s\}$ is a dominating set in $T$ and either $C_{u}\cup\{s,u\}$ or $C_{u}\cup\{l,u\}$ is an identifying code in $T$. Hence, $\gamma^{{\rm ID}}(T)\leq n-3-\gamma(T_{u})+2\leq n-\gamma(T)$ as claimed.
@@
Observe that for every tree $T$ of order at least $3$, we have $\gamma(T)\geq s(T)$ since every leaf needs to be dominated by a vertex in the dominating set. Moreover, for example, in a path, we can have $\gamma(P_{n})>s(P_{n})$. Therefore, the upper bound of Theorem 8 is an improvement over the simple but useful upper bound $n-s(T)$ from Lemma 6.
@@ 4 Characterizing extremal trees
In particular, Lemma 9 shows that $\Delta$-stars satisfy the conjectured bound with $c=\frac{1}{\Delta}$. Hence, for all stars the constant $c=\frac{1}{3}$ suffices in Conjecture 1. To fully establish the first part of Theorem 2 now, one needs to only show the veracity of the result for the rest of the trees in $\mathcal{T}_{3}$. To describe the trees in $\mathcal{T}_{3}$ and other graphs later in a more unified manner, we start by defining a particular “join” of graphs with stars. Let $G^{\prime}$ be a graph and $S$ be any star. Then, let $G^{\prime}\rhd_{v}S$ denote the graph obtained by identifying a vertex $v$ of $G^{\prime}$ with a leaf $l$ of $S$ (for example, if $S$ and $P$ are a $3$-star and $4$-path, respectively, each with a leaf $v$, then the graphs in Figures 1(a) and 1(b) are $S\rhd_{v}S$ and $P\rhd_{v}S$, respectively). We call the $G^{\prime}\rhd_{v}S$ the graph $G^{\prime}$ appended with a star and it is said to be obtained by appending $S$ (by its leaf $l$) onto (the vertex $u$ of) $G^{\prime}$. In the case that the vertex $v$ of $G^{\prime}$ is inconsequential to the context or is (up to isomorphism) immaterial to the graph $G$, we may simply drop the suffix $v$ in the notation $G^{\prime}\rhd_{v}S$ and denote it as $G^{\prime}\rhd S$ (for example, if $P$ is a $2$-path and $S$ is a star, then $P\rhd_{v}S$ is (up to isomorphism) the only graph irrespective of which vertex of the $2$-path $v$ is. Another example would be if $S$ is a $\Delta$-star and we require $S\rhd_{v}S$ to be a graph of maximum degree $\Delta$. As a convention, we continue to call the vertices of $G^{\prime}\rhd S$ by the same names as they were called in the graphs $G^{\prime}$ and $S$. In other words, the graph $G^{\prime}\rhd S$ is said to inherit its vertices from $G^{\prime}$ and $S$. In particular, if $G^{\prime}\rhd S$ is obtained by identifying the vertex $v$ of $G^{\prime}$ and a leaf $l$ of $S$, then both the names $v$ and $l$ (as and when convenient) also refer to the identified vertex in $G^{\prime}\rhd S$.
@@ 4.2 Appended stars
Case 2.2.1. $\Delta_{0}\geq 4$. We again take $h=1$ and consider the graph $G_{1}$ of order $n^{\prime}=2\Delta_{0}+1$ with $\Delta^{\prime}=\Delta_{0}\geq 4$. In this case, $G_{1}$ has $22\Delta^{\prime}-$ 2 leaves, a single degree $2$ vertex, namely $l_{1}^{2}$, and two degree $\Delta^{\prime}$ vertices, namely $u_{1}$ and $u_{2}$. The set $C^{\prime}=\{u_{1},u_{2},l_{2}^{1},\dots,l_{\Delta^{\prime}-1}^{1},l_{2}^{2},\dots,l_{\Delta^{\prime}-1}^{2}\}$ is an identifying code of $G_{1}$. Moreover, we have
@@
However, since $G_{p-1}$ has order $n-3$, this contradicts the minimality of the order of $G$. Thus, we must have $u_{i}\in C$ for all $i\in[p]$. $\blacklozenge$
@@
If ${\rm diam}(G)=4$, then in this case, $p\leq 2$ and onto exactly one leaf $c_{0}$, say, of $S_{0}$ at least one other $3$-star $S_{1}$, say, is appended by its leaf $c_{1}$. However then at least two vertices from $\{u_{0},a_{0},b_{0}\}$ must be in $C$. By the above considerations following the proof of Claim 22.A, we have at least $12p+$ 1 further vertices in $C$, and thus $|C|\geq 2p+3$. Hence, we may assume that ${\rm diam}(G)=6$, for otherwise the desired lower bound in the claim follows. In this case, $p\leq 6$ and to at least two leaves $a_{0}$ and $c_{0}$ of $S_{0}$, other $3$-stars are appended. Then again, by our previous observation, $a_{0}$ and $c_{0}$ are included in the identifying code $C$, since $a_{0}=c_{i}$ for some $i\neq 0$ and $c_{0}=c_{j}$ for some $j\neq 0$. Moreover, for $b_{0}$ to be dominated by $C$, at least one of $\{u_{0},b_{0}\}$ must be in $C$. Thus, we again have $|C|\geq 2p+3$. This proves the claim. $\blacklozenge$
@@
Let $G^{\prime}=G-\{c_{0}\}$. Observe that $G^{\prime}$ consists of two identical components of order $6$. We show that each of them has at least four vertices in the identifying code. First of all, $a_{0}\in C$ is forced as the only vertex which separates $u_{0}$ and $b_{0}$. Moreover, |$|\{u_{0},b_{0}\}\cap C|\geq 1$ since $C$ dominates $b_{0}$. Finally, we have |$|\{a_{2},u_{2},b_{2}\}\cap C|\geq 2$ to dominate and separate $a_{2}$ and $b_{2}$. Thus, $|C|\geq 8$. $\blacklozenge$
@@
    If $T\notin\{T_{2},T_{3}\}$, then $C(T)$ can be chosen as an independent set which contains every vertex of degree at most $2$. When we delete any code vertex $v$ from $T$, set $C(T)\setminus\{v\}$ forms an optimal identifying code of the forest $T-v$.
@@ 6 Tightness of the bound for trees
Moreover, it is tight for some infinite families (with $c=0$), such as double stars with $22\Delta-$ 2 leaves, where by a double star we mean a tree $G=S_{1}\rhd_{u}S_{2}$, where $S_{1}$ is a $(\Delta-1)$-star, $S_{2}$ is a $\Delta$-star and $u$ is universal vertex of $S_{1}$. The double star $G$ has $n=2\Delta$ vertices and the smallest identifying code $C$ of $G$ has $22\Delta-$ 2 code vertices. Thus, $|C|=(\frac{\Delta-1}{\Delta})n$.
@@
$\frac{\Delta-1+\frac{1}{\Delta-2}}{\Delta+\frac{2}{\Delta-2}}=$ $\frac{1}{\Delta}\left(\frac{\Delta^{2}-\Delta+\frac{\Delta}{\Delta-2}}{\Delta+\frac{2}{\Delta-2}}\right)\vspace{0.2cm}$
$=$ $\frac{1}{\Delta}\left(\frac{\Delta^{3}-3\Delta^{2}+3\Delta}{\Delta^{2}-2\Delta+2}\right)\vspace{0.2cm}$
$=$ $\frac{\Delta-1}{\Delta}-\frac{\Delta-2}{\Delta^{3}-2\Delta^{2}+2\Delta}\vspace{0.2cm}$
$>$ $\frac{\Delta-1}{\Delta}-\frac{\Delta-2}{\Delta^{2}(\Delta-2)}\vspace{0.2cm}$

### v3

@@ 1 Introduction
A natural question that arises in the study of identifying codes (like for the usual dominating sets) is the one of extremal values for the identification number: how large can it be, with respect to some relevant graph parameters? When only the order $n$ of the graph is considered, it is known that the identification number of an identifiable graph with at least one edge lies between $\log_{2}(n+1)$ [33] and n1n-1 [23]; both values are tight and the extremal examples have been characterized in [36] and [15], respectively.
@@
With respect to the identification number, it was observed in [15] that when the maximum degree $\Delta$ of the graph $G$ is small enough with respect to the order $n$ of the graph, the (n1n-1)-upper bound can be significantly improved (for connected graphs) to $n-\frac{n}{\Theta(\Delta^{5})}$. The latter was thereafter subsequently reduced to $n-\frac{n}{\Theta(\Delta^{3})}$ in [18]. This raises the question of what is the largest possible identification number of a connected identifiable graph of order $n$ and maximum degree $\Delta$. Towards this problem, the following conjecture was posed; the goal of this paper is to study this conjecture.
@@
Even paths 2 11 [5] (Theorem 5)
@@ 2.1 Definitions and notations
For any vertex $v$ of $G$, the symbol $\deg_{G}(v)$ denoting the degree of the vertex $v$ in $G$ is the total number of neighbors of $v$ in $G$. A leaf of a graph $G$ is a vertex of degree $1$ in $G$. In the literature, a leaf is also known as a pendant vertex. The (only) neighbor of a leaf $v$ in a graph $G$ is called the support vertex of $v$ in $G$. The number of leaves and support vertices in $G$ are denoted by $\ell(G)$ and $s(G)$, respectively. Naturally, any vertex of a graph $G$ that is not a leaf of $G$ is usually referred to as a non-leaf vertex of $G$. The length (or the number of edges) of a longest induced path in a graph $G$ is called the diameter of $G$ and is denoted by ${\rm diam}(G)$.
On many occasions throughout this article, we shall have the need to look at a subgraph of a graph $G$ formed by deleting away some vertices or edges from $G$. To that end, given a graph $G$ and a set $S$ containing some vertices and edges of $G$, we define GSG-S as the subgraph of $G$ obtained by deleting from $G$ all vertices (and edges incident with them) and edges of $G$ in $S$. For any positive integer $p$, let $[p]$ denote the set $\{1,2,\ldots,p\}$ and for any two distinct integers $p$ and $q$ with $p<q$, let $[p,q]$ denote the set $\{p,p+1,\ldots,q\}$.
@@ 2.2 Paths
$\gamma^{{\rm ID}}(P_{n})=\begin{cases}\frac{n}{2}+\frac{1}{2},&\text{if $n\geq 1$ is odd},\vskip 2.84544pt\\ \frac{n}{2}+1,&\text{if $n\geq 4$ is even}.\end{cases}$
@@ 3 An upper bound relating the identification number of a tree with its domination number
Let us first assume that $s$ is adjacent to $t\geq 2$ leaves $l_{1},l_{2},\dots,l_{t}$. Consider the tree $T_{s}=T-\{s,l_{1},l_{2},\dots,l_{t}\}$. Since $T_{s}$ has nt1n-t-1 vertices, by induction we have $\gamma^{{\rm ID}}(T_{s})\leq n-t-1-\gamma(T_{s})$ unless $T_{s}=P_{4}$, or $T_{s}$ contains at most two vertices. However, notice that $|V(T_{s})|\geq 3$ on account of $x$ being a non-leaf and non-support vertex of $T$. Moreover, if $T_{s}=P_{4}$, then $s$ cannot be adjacent to a support vertex of $T_{s}$, or else, the vertex $x$ would be either a leaf or a support vertex of $T$, a contradiction to our assumption. Therefore the vertex $s$ must be adjacent to a leaf of $T_{s}$. In this case, we have $\gamma(T)=2=s(T)$. Since $\gamma^{{\rm ID}}(T)\leq n-s(T)$ by Lemma 6, the claim holds in this case. Therefore, we may assume by induction that there exists an identifying code $C_{s}$ in $T_{s}$ which has cardinality at most $n-t-1-\gamma(T_{s})$, with $D_{s}$ a minimum-size dominating set in $T_{s}$. Furthermore, set $C_{s}\cup\{l_{1},\dots,l_{t}\}$ is an identifying code of $T$, and set $D_{s}\cup\{s\}$ is a dominating set of $T$. Thus, $\gamma^{{\rm ID}}(T)\leq n-t-1-\gamma(T_{s})+t\leq n-\gamma(T)$, as claimed.
Let us assume next that support vertex $s$ has degree 2, and denote by $u$ the non-leaf adjacent to $s$ (possibly, $u=x$) and by $l$ the leaf adjacent to $s$. If vertex $u$ also has degree 2, then we consider the tree $T_{u}=T-\{l,s,u\}$. Since $x$ is a non-leaf and non-support vertex of $T$, we must have $|V(T_{u})|\geq 2$. However, if $|V(T_{u})|=2$, then $T$ is a path, a contradiction. Moreover, if $T_{u}\cong P_{4}$, then $T$ is path if $u$ is adjacent to a leaf of $T_{u}$, again a contradiction. Hence, $T$ has three support vertices and $\gamma^{{\rm ID}}(T)\leq n-3$ (consider the identifying code formed by all non-leaf vertices of $T$) while $\gamma(T)=3$. Hence, by induction, we may assume that $T_{u}$ has an optimal identifying code $C_{u}$ with cardinality $|C_{u}|=\gamma^{{\rm ID}}(T_{u})\leq n-3-\gamma(T_{u})$ and we consider a minimum-size dominating set $D_{u}$ of $T_{u}$. Notice that $D_{u}\cup\{s\}$ is a dominating set in $T$ and either $C_{u}\cup\{s,u\}$ or $C_{u}\cup\{l,u\}$ is an identifying code in $T$. Hence, $\gamma^{{\rm ID}}(T)\leq n-3-\gamma(T_{u})+2\leq n-\gamma(T)$ as claimed.
@@
Observe that for every tree $T$ of order at least $3$, we have $\gamma(T)\geq s(T)$ since every leaf needs to be dominated by a vertex in the dominating set. Moreover, for example, in a path, we can have $\gamma(P_{n})>s(P_{n})$. Therefore, the upper bound of Theorem 8 is an improvement over the simple but useful upper bound ns(T)n-s(T) from Lemma 6.
@@ 4 Characterizing extremal trees
In particular, Lemma 9 shows that $\Delta$-stars satisfy the conjectured bound with $c=\frac{1}{\Delta}$. Hence, for all stars the constant $c=\frac{1}{3}$ suffices in Conjecture 1. To fully establish the first part of Theorem 2 now, one needs to only show the veracity of the result for the rest of the trees in $\mathcal{T}_{3}$. To describe the trees in $\mathcal{T}_{3}$ and other graphs later in a more unified manner, we start by defining a particular “join” of graphs with stars. Let $G^{\prime}$ be a graph and $S$ be any star. Then, let $G^{\prime}\rhd_{v}S$ denote the graph obtained by identifying a vertex $v$ of $G^{\prime}$ with a leaf $l$ of $S$ (for example, if $S$ and $P$ are a $3$-star and $4$-path, respectively, each with a leaf $v$, then the graphs in Figures 1(b) and 1(c) are $S\rhd_{v}S$ and $P\rhd_{v}S$, respectively). We call the $G^{\prime}\rhd_{v}S$ the graph $G^{\prime}$ appended with a star and it is said to be obtained by appending $S$ (by its leaf $l$) onto (the vertex $v$ of) $G^{\prime}$. In the case that the vertex $v$ of $G^{\prime}$ is inconsequential to the context or is (up to isomorphism) immaterial to the graph $G$, we may simply drop the suffix $v$ in the notation $G^{\prime}\rhd_{v}S$ and denote it as $G^{\prime}\rhd S$ (for example, if $P$ is a $2$-path and $S$ is a star, then $P\rhd_{v}S$ is (up to isomorphism) the only graph irrespective of which vertex of the $2$-path $v$ is. Another example would be if $S$ is a $\Delta$-star and we require $S\rhd_{v}S$ to be a graph of maximum degree $\Delta$. As a convention, we continue to call the vertices of $G^{\prime}\rhd S$ by the same names as they were called in the graphs $G^{\prime}$ and $S$. In other words, the graph $G^{\prime}\rhd S$ is said to inherit its vertices from $G^{\prime}$ and $S$. In particular, if $G^{\prime}\rhd S$ is obtained by identifying the vertex $v$ of $G^{\prime}$ and a leaf $l$ of $S$, then both the names $v$ and $l$ (as and when convenient) also refer to the identified vertex in $G^{\prime}\rhd S$.
@@ 4.2 Appended stars
Case 2.2.1. $\Delta_{0}\geq 4$. We again take $h=1$ and consider the graph $G_{1}$ of order $n^{\prime}=2\Delta_{0}+1$ with $\Delta^{\prime}=\Delta_{0}\geq 4$. In this case, $G_{1}$ has $2\Delta^{\prime}-2$ leaves, a single degree $2$ vertex, namely $l_{1}^{2}$, and two degree $\Delta^{\prime}$ vertices, namely $u_{1}$ and $u_{2}$. The set $C^{\prime}=\{u_{1},u_{2},l_{2}^{1},\dots,l_{\Delta^{\prime}-1}^{1},l_{2}^{2},\dots,l_{\Delta^{\prime}-1}^{2}\}$ is an identifying code of $G_{1}$. Moreover, we have
@@
However, since $G_{p-1}$ has order n3n-3, this contradicts the minimality of the order of $G$. Thus, we must have $u_{i}\in C$ for all $i\in[p]$. $\blacklozenge$
@@
If ${\rm diam}(G)=4$, then in this case, $p\leq 2$ and onto exactly one leaf $c_{0}$, say, of $S_{0}$ at least one other $3$-star $S_{1}$, say, is appended by its leaf $c_{1}$. However then at least two vertices from $\{u_{0},a_{0},b_{0}\}$ must be in $C$. By the above considerations following the proof of Claim 22.A, we have at least $2p+1$ further vertices in $C$, and thus $|C|\geq 2p+3$. Hence, we may assume that ${\rm diam}(G)=6$, for otherwise the desired lower bound in the claim follows. In this case, $p\leq 6$ and to at least two leaves $a_{0}$ and $c_{0}$ of $S_{0}$, other $3$-stars are appended. Then again, by our previous observation, $a_{0}$ and $c_{0}$ are included in the identifying code $C$, since $a_{0}=c_{i}$ for some $i\neq 0$ and $c_{0}=c_{j}$ for some $j\neq 0$. Moreover, for $b_{0}$ to be dominated by $C$, at least one of $\{u_{0},b_{0}\}$ must be in $C$. Thus, we again have $|C|\geq 2p+3$. This proves the claim. $\blacklozenge$
@@
Let $G^{\prime}=G-\{c_{0}\}$. Observe that $G^{\prime}$ consists of two identical components of order $6$. We show that each of them has at least four vertices in the identifying code. First of all, $a_{0}\in C$ is forced as the only vertex which separates $u_{0}$ and $b_{0}$. Moreover, $|\{u_{0},b_{0}\}\cap C|\geq 1$ since $C$ dominates $b_{0}$. Finally, we have $|\{a_{2},u_{2},b_{2}\}\cap C|\geq 2$ to dominate and separate $a_{2}$ and $b_{2}$. Thus, $|C|\geq 8$. $\blacklozenge$
@@
    If $T\notin\{T_{2},T_{3}\}$, then $C(T)$ can be chosen as an independent set which contains every vertex of degree at most $2$. When we delete any code vertex $v$ from $T$, set $C(T)\setminus\{v\}$ forms an optimal identifying code of the forest TvT-v.
@@ 6 Tightness of the bound for trees
Moreover, it is tight for some infinite families (with $c=0$), such as double stars with $2\Delta-2$ leaves, where by a double star we mean a tree $G=S_{1}\rhd_{u}S_{2}$, where $S_{1}$ is a $(\Delta-1)$-star, $S_{2}$ is a $\Delta$-star and $u$ is universal vertex of $S_{1}$. The double star $G$ has $n=2\Delta$ vertices and the smallest identifying code $C$ of $G$ has $2\Delta-2$ code vertices. Thus, $|C|=(\frac{\Delta-1}{\Delta})n$.
@@
$\frac{\Delta-1+\frac{1}{\Delta-2}}{\Delta+\frac{2}{\Delta-2}}=$ $\frac{1}{\Delta}\left(\frac{\Delta^{2}-\Delta+\frac{\Delta}{\Delta-2}}{\Delta+\frac{2}{\Delta-2}}\right)\vskip 5.69046pt$
$=$ $\frac{1}{\Delta}\left(\frac{\Delta^{3}-3\Delta^{2}+3\Delta}{\Delta^{2}-2\Delta+2}\right)\vskip 5.69046pt$
$=$ $\frac{\Delta-1}{\Delta}-\frac{\Delta-2}{\Delta^{3}-2\Delta^{2}+2\Delta}\vskip 5.69046pt$
$>$ $\frac{\Delta-1}{\Delta}-\frac{\Delta-2}{\Delta^{2}(\Delta-2)}\vskip 5.69046pt$